OpenStudy (anonymous):

Find the inverse laplace transform of alpha/((s^2+2s+2)(s^2+(alpha)^2)?

OpenStudy (loser66):

\[\frac{\alpha}{s^2+\alpha^2}~*~\frac{1}{(s^2+2s+1)+1}\\\frac{\alpha}{s^2+\alpha^2}~*~\frac{1}{(s+1)^2+1}\] \[L^{-1}\left[\frac{\alpha}{s^2+\alpha^2}\right]=sin \alpha t\] \[L^{-1}\left[\frac{1}{(s+1)^2 +1}\right]=e^{-t}sin~t\] now, combine. take G(s) as the first one, and F(s) as the second one. You have the form of F(s)*G(s) formula #16 tells you that \[\int_o^t f(t-\tau)g(\tau)d\tau\]

OpenStudy (anonymous):

Sorry, I don't understand what you wrote. The processing math is 0%.

OpenStudy (loser66):